# HBSplat: Robust Sparse-View Gaussian Reconstruction with Hybrid-Loss Guided Depth and Bidirectional Warping

### v2

@@ Abstract
(a) Under Sparse-View Reconstruction: Rendered images and depth maps comparison between MCGS and HBSplat.
@@
Figure 1: (a) MCGS outputs (left) vs. HBSplat outputs (right) for both image and depth map. (b) presents an efficiency-quality scatter plot comparing HBSplat with various baseline methods.
@@ I Introduction
Figure 2: HBSplat pipeline. First, sparse input images are processed by dense matching, SfM, and monocular depth estimation to obtain correspondences, camera poses, and depth maps. The Hybrid-Loss module fuses these inputs to produce robust point-wise depths. Subsequent least-squares optimization aligns the point cloud with monocular depths, recovering the metric scale. The Bidirectional Warping module leverages these aligned depths and images to synthesize novel virtual training views through depth-image warping and interpolation. Simultaneously, the Occlusion-Aware Reconstruction component restores missing background content in occluded regions using learning-based inpainting model (Simple-LAMA) guided by local foreground mask. Finally, the framework reconstructs the 3D Gaussian scene by optimizing a joint loss that combines color and depth supervision.
@@ II-1 NeRF-Based NVS Methods
Neural Radiance Fields (NeRF) [4] implicitly represent scenes via Multi-Layer Perceptrons (MLPs), achieving photorealistic NVS through volume rendering and becoming a dominant approach in recent years. Subsequent improvements include Mip-NeRF [13], which mitigates aliasing via conical frustum rendering; Mip-NeRF360 [14], extending the framework to 360° unbounded scenes; and InstantNGP [15], which reduces training time to seconds using multi-resolution hash encoding. However, these methods generally require dense view inputs (often dozens of images) and suffer from slow training and inference. Moreover, in sparse-view settings, NeRF-based approaches are prone to overfitting, resulting in blurry or inconsistent novel views.
@@
3DGS represents scenes with explicit Gaussian primitives, achieving real-time rendering via differentiable rasterization, surpassing NeRF in efficiency and quality. 3DGS excels with dense views, supporting tasks like text-to-3D generation [23] and dynamic scene modeling [24, 25]. However, in sparse views, 3DGS suffers from insufficient initial point clouds and limited multi-view constraints, leading to floating artifacts and geometry degradation.
@@ III-C1 Initial Depth Estimation Stage
Reprojection Constraint ($RC$): Using dense image matching correspondences, such as point pair $a$-$b$ between image pair II-$J$, point $a$ from image $I$ is back-projected into 3D space to obtain point $A$, which is then projected into image $J$ to yield $a^{\prime}$. The reprojection loss minimizes the Euclidean distance between the reprojected and original points $a^{\prime}$-$b$:
@@
Point Propagation Constraint ($PPC$): Unlike sparse matcher that independently extract feature points, dense matcher generate matching points based on image pairs, making reprojection error constraints alone insufficient for ensuring depth consistency and robustness. To address this, the Point Propagation Constraint is introduced: for matching point pair $a$-$b$ between image II-$J$, and $b^{\prime}$-$c$ between image pair $J$-$K$, we use point $b$ and $b^{\prime}$ in image $J$ as a bridge to establish indirect geometric consistency between $a$ and $c$ across views. The common point set $\mathcal{C}$ is defined as:
@@
where $\mathcal{M}_{i,j}$ denotes matched point pairs between images $I$ and $J$, and $d_{NN}$ is a mutual nearest-neighbor distance threshold. Since identical physical points may not align exactly in pixel coordinates, a mutual nearest-neighbor search is performed in a KD-Tree: for each $p_{j}^{b}$, its nearest neighbor $p_{j}^{b^{\prime}}$ is identified, and vice versa. If the distance between them is less than $d_{\text{NN}}$, they are considered the same point. The point propagation loss, similar to reprojection loss is formulated as:
@@
0: Matching point pairs $\mathcal{M}_{i,j}={(\mathbf{p}_{i}^{a},\mathbf{p}_{j}^{b})}$ between views $i$ and $j$, $\mathcal{M}_{j,k}={(\mathbf{p}_{j}^{b^{\prime}},\mathbf{p}_{k}^{c})}$ between views $j$ and $k$, pixel distance threshold $\epsilon$;
@@
9:      if $\mathbf{p}_{j}^{b}=\mathbf{p}_{j}^{b^{\prime\prime}}$ and $|\mathbf{p}_{j}^{b}-\mathbf{p}_{j}^{b^{\prime}}|<\epsilon$ then
@@
$\begin{cases}\tau_{\text{dy}}=\tau_{\text{base}}+\alpha\cdot Sigmoid(2\hat{z}-1),\\[5.69054pt] M_{i,j}^{depth}=\frac{|z_{ij}-z_{ji}|}{\min(z_{ij},z_{ji})}\leq\tau_{\text{dy}},\\[5.69054pt] M_{ij}^{error}=E_{ij}^{RC}\leq\tau_{RC},\\[5.69054pt] M_{i,j}^{RC\text{-}dep}=M_{ij}^{error}\odot M_{ji}^{error}\odot M_{i,j}^{depth},\end{cases}$ (9)
where the dynamic threshold $\tau_{\text{dy}}$ is implemented using a Sigmoid function to enforce stricter depth consistency constraints for nearby points and more relaxed tolerances for distant points; $\tau_{\text{base}}$ denotes the base threshold, $\alpha$ is a scaling factor, and $\hat{z}$ represents the normalized depth value; $z_{ij}$ denotes the ray depth of image $I$ in the image pair II-$J$; $M_{i,j}^{\text{depth}}$ is the depth consistency mask, obtained by comparing the relative depth difference with $\tau_{\text{dy}}$; $E_{ij}^{RC}$ denotes the reprojection error between images $I$ and $J$; $\tau_{RC}$ is a fixed reprojection error threshold; $M_{ij}^{\text{error}}$ is the reprojection error mask; $M_{i,j}^{RC\text{-}dep}$ is the final valid matching mask that combines bidirectional reprojection errors and depth consistency. Based on this mechanism, matched point pairs are filtered if either the reprojection error exceeds the predefined threshold or the relative depth difference is too large.
@@ III-C2 Scene Training Stage
Smoothness Constraint: Total Variation (TV) loss is applied to promote spatial smoothness in the rendered depth map.
@@ IV-A4 Implementation Details
HBSplat is built upon the 3DGS framework[5]. Camera poses are estimated using COLMAP across all views for robustness. To ensure consistent comparison, similar to SCGaussian, we employed the GIM-DKM[44] model for dense matching to extract matching points per training view. Besides, we use ML-depth-pro [45] for monocular depth estimation, which excels in capturing fine details. Both depth estimation and scene training are iterated 2,000 times. Initialize and generate Gaussian primitives after filtering outliers. Bidirectional Warping generates 50 virtual views; one virtual view iteration is performed for every 5 real view iterations. The Occlusion-Aware Reconstruction applies a depth difference mask to prioritize rendering of occluded region Gaussians in the first 1,000 iterations, followed by training all Gaussians. Gaussian cloning and pruning follow the vanilla 3DGS settings. All experiments are conducted on a single NVIDIA RTX 3090 GPU. Please refer to the publicly available code for detailed configuration.
@@
Method 1/8 LLFF 1/4 LLFF 1/8 IBRNet
@@ IV-B5 Efficiency
Efficiency evaluation of different methods on the LLFF dataset in Table VII. The proposed HBSplat method matches the efficiency of other advanced approaches, achieving an inference speed of 250 FPS, which is comparable to methods such as SCGaussian and MCGaussian, while maintaining a low average training time of 2.5 minutes. In terms of GPU memory usage, HBSplat optimizes resource consumption, requiring 3 GB for the Hybrid-Loss Depth Estimation (HLDE) component and 5 GB for the Virtual View Synthesis (VVS) component, outperforming methods like FreeNeRF (4×48 GB) and SparseNeRF (32 GB). Compared to approaches such as 3DGS (7.5 minutes, 2 GB) and FSGS (17 minutes, 3 GB), HBSplat offers a substantial reduction in training time while preserving real-time inference capabilities, making it a highly efficient solution for practical 3D reconstruction tasks.
TABLE VII: Efficiency comparison of different methods.
@@ IV-C Ablation Study
In the second row of the comparison, SCGaussian exhibits black artifacts and fails to reconstruct the background completely, while MCGS reconstructs leaves as sparse, needle-like structures and suffers from blurred edges. In contrast, HBSplat produces renderings that are visually closer to the GT, demonstrating the effectiveness of the proposed components.
@@
Figure 13: Ablation study on component contributions. Compared to MCGS and SCGaussian, the rendering results of full HBSplat model are closest to.
@@ V Conclusion
This paper presents HBSplat, an effective framework for sparse-view NVS based on 3DGS. The core of our approach includes three innovations: a Hybrid-Loss Depth Estimation module that enforces multi-view consistency through reprojection, point propagation, and smoothness constraints; a Bidirectional Warping Virtual View Synthesis method that enhances the coverage of unobserved regions, thereby mitigating overfitting to the limited input views; and an Occlusion-Aware Reconstruction component that improves background rendering through depth-difference priors. Extensive experiments across standard benchmarks demonstrate that HBSplat achieves state-of-the-art performance under extreme sparsity. We believe our work offers a valuable step toward practical and high-quality 3D reconstruction from very few images.

### v3

@@ Abstract
(a) Under Sparse-View Reconstruction: Rendered images and depth maps comparison between MCGS (left) and HBSplat (right).
@@
Figure 1: (a) MCGS outputs (left) vs. HBSplat outputs (right) for both rendered image and depth map. (b) presents an efficiency-quality scatter plot comparing HBSplat with various baseline methods.
@@ I Introduction
Figure 2: HBSplat pipeline. First, sparse input images are processed by dense matching, structure from motion (SfM), and monocular depth estimation to obtain correspondences, camera poses, and depth maps. The Hybrid-Loss module fuses these inputs to produce robust point-wise depths. Subsequent least-squares optimization aligns the point cloud with monocular depths, recovering the metric scale. The Bidirectional Warping module leverages these aligned depths and images to synthesize novel virtual training views through depth-image warping and interpolation. Simultaneously, the Occlusion-Aware Reconstruction component restores missing background content in occluded regions using learning-based inpainting model (Simple-LAMA) guided by local foreground mask. Finally, the framework reconstructs the 3D Gaussian scene by optimizing a joint loss that combines color and depth supervision.
@@ II-1 NeRF-Based NVS Methods
NeRF [4] implicitly represent scenes via Multi-Layer Perceptrons (MLPs), achieving photorealistic NVS through volume rendering and becoming a dominant approach in recent years. Subsequent improvements include Mip-NeRF [13], which mitigates aliasing via conical frustum rendering; Mip-NeRF360 [14], extending the framework to 360° unbounded scenes; and InstantNGP [15], which reduces training time to seconds using multi-resolution hash encoding. However, these methods generally require dense view inputs (often dozens of images) and suffer from slow training and inference. Moreover, in sparse-view settings, NeRF-based approaches are prone to overfitting, resulting in blurry or inconsistent novel views.
@@
3DGS represents scenes with explicit Gaussian primitives, achieving real-time rendering via differentiable rasterization, surpassing NeRF in efficiency and quality. 3DGS excels with dense views, supporting tasks like text-to-3D generation [23] and dynamic scene modeling [24, 25]. However, in sparse views, 3DGS suffers from insufficient initial point clouds and limited multi-view constraints, leading to artifacts and geometry degradation.
@@ III-C1 Initial Depth Estimation Stage
Reprojection Constraint ($RC$): Using dense image matching correspondences, such as point pair $a$-$b$ between view pair $i$-$j$, point $a$ from view $i$ is back-projected into 3D space to obtain point $A$, which is then projected into view $j$ to yield $a^{\prime}$. The reprojection loss minimizes the Euclidean distance between the reprojected and original points $a^{\prime}$-$b$:
@@
Point Propagation Constraint ($PPC$): Unlike sparse matcher that independently extract feature points, dense matcher generate matching points based on image pairs, making reprojection error constraints alone insufficient for ensuring depth consistency and robustness. To address this, the Point Propagation Constraint is introduced: for matching point pair $a$-$b$ between view pair $i$-$j$, and $b^{\prime}$-$c$ between view pair $i$-$k$, we use point $b$ and $b^{\prime}$ in view $j$ as a bridge to establish indirect geometric consistency between $a$ and $c$ across views. The common point set $\mathcal{C}$ is defined as:
@@
where $\mathcal{M}_{i,j}$ denotes matched point pairs between views $i$ and $j$, and $d_{NN}$ is a mutual nearest-neighbor distance threshold. Since identical physical points may not align exactly in pixel coordinates, a mutual nearest-neighbor search is performed in a KD-Tree: for each $p_{j}^{b}$, its nearest neighbor $p_{j}^{b^{\prime}}$ is identified, and vice versa. If the distance between them is less than $d_{\text{NN}}$, they are considered the same point. The point propagation loss, similar to reprojection loss is formulated as:
@@
0: Matching point pairs $\mathcal{M}_{i,j}={(\mathbf{p}_{i}^{a},\mathbf{p}_{j}^{b})}$ between views $i$ and $j$, $\mathcal{M}_{j,k}={(\mathbf{p}_{j}^{b^{\prime}},\mathbf{p}_{k}^{c})}$ between views $j$ and $k$, pixel distance threshold $d_{NN}$;
@@
9:      if $\mathbf{p}_{j}^{b}=\mathbf{p}_{j}^{b^{\prime\prime}}$ and $|\mathbf{p}_{j}^{b}-\mathbf{p}_{j}^{b^{\prime}}|<d_{NN}$ then
@@
$\begin{cases}\tau_{dy}=\tau_{base}+\alpha\cdot Sigmoid(2\hat{z}-1),\\[5.69054pt] M_{i,j}^{depth}\leftarrow\mathcal{D}(z_{ij},z_{ji})=\frac{|z_{ij}-z_{ji}|}{\min(z_{ij},z_{ji})}\leq\tau_{dy},\\[5.69054pt] M_{ij}^{error}=E_{ij}^{RC}\leq\tau_{RC},\\[5.69054pt] M_{i,j}^{RC\text{-}dep}=M_{ij}^{error}\odot M_{ji}^{error}\odot M_{i,j}^{depth},\end{cases}$ (9)
where the dynamic threshold $\tau_{dy}$ is implemented using a Sigmoid function to enforce stricter depth consistency constraints for nearby points and more relaxed tolerances for distant points; $\tau_{base}$ denotes the base threshold, $\alpha$ is a scaling factor, and $\hat{z}$ represents the normalized depth value; $z_{ij}$ denotes the ray depth of view $i$ in the view pair $i$-$j$; $M_{i,j}^{depth}$ is the depth consistency mask, obtained by comparing the relative depth difference with $\tau_{dy}$; $E_{ij}^{RC}$ denotes the reprojection error between views $i$ and $j$; $\tau_{RC}$ is a fixed reprojection error threshold; $M_{ij}^{error}$ is the reprojection error mask; $M_{i,j}^{RC\text{-}dep}$ is the final valid matching mask that combines bidirectional reprojection errors and depth consistency. Based on this mechanism, matched point pairs are filtered if either the reprojection error exceeds the predefined threshold or the relative depth difference is too large.
@@ III-C2 Scene Training Stage
Smoothness Constraint: The Total Variation (TV) loss is applied to promote spatial smoothness in the rendered depth map:
@@ IV-A4 Implementation Details
HBSplat is built upon the 3DGS framework[5]. Camera poses are estimated using COLMAP across all views for robustness. To ensure consistent comparison, similar to SCGaussian, we employed the GIM-DKM[44] model for dense matching to extract matching points per training view. Besides, we use ML-depth-pro [45] for monocular depth estimation, which excels in capturing fine details. Both depth estimation and scene training are iterated 2,000 times. Initialize and generate Gaussian primitives after filtering outliers. Bidirectional Warping generates 100 virtual views; one virtual view iteration is performed for every 5 real view iterations. The Occlusion-Aware Reconstruction applies a depth difference mask to prioritize rendering of occluded region Gaussians in the first 1,000 iterations, followed by training all Gaussians. Gaussian cloning and pruning follow the vanilla 3DGS settings. All experiments are conducted on a single NVIDIA RTX 3090 GPU. For detailed configuration, please refer to the publicly available code.
@@
Method 1/8 LLFF 1/4 LLFF 1/2 IBRNet
@@ IV-B5 Efficiency
Efficiency evaluation of different methods on the LLFF dataset in Table VII. The proposed HBSplat method matches the efficiency of other advanced approaches, achieving an inference speed of 250 FPS, which is comparable to methods such as SCGaussian and MCGaussian, while maintaining a low average training time of 2.5 minutes. For GPU memory, HBSplat requires 3 GB for training and an additional 5 GB for the Virtual View Synthesis (VVS) component to generate 100 virtual views, outperforming methods like FreeNeRF (4×48 GB) and SparseNeRF (32 GB). Compared to approaches such as 3DGS (7.5 minutes, 2 GB) and FSGS (17 minutes, 3 GB), HBSplat offers a substantial reduction in training time while preserving real-time inference capabilities, making it a highly efficient solution for practical 3D reconstruction tasks.
TABLE VII: Efficiency comparison of different methods on LLFF dataset.
@@ IV-C Ablation Study
In the second row of the comparison, MCGS reconstructs leaves as sparse, needle-like structures and suffers from blurred edges. SCGaussian exhibits black artifacts and fails to reconstruct the background completely. In contrast, HBSplat produces renderings that are visually closer to the GT, demonstrating the effectiveness of the proposed components.
@@
Figure 13: Ablation study on component contributions. Compared to MCGS and SCGaussian, the rendering results of full HBSplat model are closest to GT.
@@ V Conclusion
This paper presents HBSplat, an effective framework for sparse-view NVS based on 3DGS. The core of our approach includes three innovations: a Hybrid-Loss Depth Estimation module that enforces multi-view consistency through reprojection, point propagation, and TV smoothness constraints; a Bidirectional Warping Virtual View Synthesis method that enhances the coverage of unobserved regions, thereby mitigating overfitting to the limited input views; and an Occlusion-Aware Reconstruction component that improves background rendering through depth-difference priors. Extensive experiments across standard benchmarks demonstrate that HBSplat achieves state-of-the-art performance under extreme sparsity. We believe our work offers a valuable step toward practical and high-quality 3D reconstruction from very few images.